### v2

# EA-KD: Entropy-based Adaptive Knowledge Distillation

Abstract
Knowledge distillation (KD) enables a smaller “student” model to mimic a larger “teacher” model by transferring knowledge from the teacher’s output or features. However, most KD methods treat all samples uniformly, overlooking the varying learning value of each sample and thereby limiting effectiveness. In this paper, we propose Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that prioritizes learning from valuable samples. EA-KD quantifies each sample’s learning value by strategically combining the entropy of the teacher and student output, then dynamically reweights the distillation loss to place greater emphasis on high-value samples. Extensive experiments across diverse KD frameworks and tasks–including image classification, object detection, and large language model (LLM) distillation–demonstrate that EA-KD consistently enhances performance, achieving state-of-the-art results with negligible computational cost. Our code will be publicly available.
11footnotetext: Equal contribution22footnotetext: Corresponding author: camhero@gmail.com
1 Introduction
The increasing size of state-of-the-art (SOTA) deep learning models in computer vision (CV) [32, 21, 6], and large language models (LLMs) [5, 25] pose challenges for deployment in resource-constrained settings. Knowledge Distillation (KD) [9] offers a solution by training a smaller “student” model to mimic a larger “teacher” model, using both ground-truth labels and the teacher’s “knowledge” (e.g. softened logits or feature representations) to achieve similar performance in a compact form.
Figure 1: Training Time per Batch (ms) vs. Testing Accuracy (%) on CIFAR-100. Highlighting EA methods’ enhanced performance with minimal additional training cost.
SOTA KD methods have explored diverse knowledge forms and structural modifications to refine knowledge transfer [9, 26, 1, 24, 38, 11, 19]. However, most methods distill uniformly across all samples, operating under the assumption that each sample has equal importance and overlooking their varying learning needs. Recent advancements, such as Instance-T [16], have shown that assigning unique temperatures to each sample outperforms the uniform temperature method Global-T. This highlighted the benefits of adapting to each sample’s distinct characteristics. Building on this, we hypothesize that emphasizing samples rich in valuable knowledge [^1] Throughout this paper, we consider samples with high entropy as containing valuable knowledge. can further optimize the distillation process. This mimics how human students learn better when key points are highlighted by teachers, and thus, this targeted approach allows the student model to achieve better comprehension by focusing on more informative samples. As we will demonstrate, the uniform distillation in most KD methods also tends to overshadow these important samples, limiting the efficiency of knowledge transfer.
The question then arises: How can we identify the most valuable samples for learning? Entropy, the core concept in information theory that quantifies the uncertainty or information of a random variable [28], may be well-suited for this role. While a prior multi-teacher KD study utilized inverse entropy to measure the reliability of teachers’ output [13], this method does not apply to the more prevalent single-teacher setting. Moreover, relying solely on the teacher’s entropy restricts the adaptability to the evolving learning dynamics of the student, leaving the full potential of entropy in KD under-explored. Therefore, we propose leveraging entropy to quantify the learning value of each sample in KD, as high-entropy outputs shall correspond to greater information contents that are crucial for learning. Our preliminary analysis reveals that higher entropy samples [^2] Results for high teacher entropy samples are shown here; high student entropy plots are in the Sec. A.1. (i) correlate with larger teacher-student accuracy gaps (Fig. 2(a)) and (ii) often lie near class boundaries in t-SNE visualizations [33] (Fig. 2(b)), suggesting these informative samples not only offer valuable learning opportunities but are also pivotal in defining decision boundaries. Hence, entropy can serve as a reliable metric for identifying valuable samples that are crucial for learning in KD.
(a) Accuracy vs. Entropy Segments.
(b) t-SNE of Teacher.
Figure 2: High-Entropy Samples in KD. (a) Higher teacher entropy correlates with larger accuracy gaps in KD, while EA-KD maintains closer alignment. (b) The top 10% of high entropy samples (denoted by triangles) cluster near decision boundaries, representing critical knowledge essential for classification.
Table 1: Entropy-based Reweighting on KD. Reweighting with entropy ($H^{\mathcal{T}}$ and $H^{\mathcal{S}}$) improves student’s performance, while inverted reweighting ($H_{\text{ub}}-w$) reduces accuracy.
Reweighting Factor $w$
Loss Function $H^{\mathcal{T}}$ $H^{\mathcal{S}}$
$L_{\text{KD}}$ [9] 73.33
$wL_{\text{KD}}$ 75.14 +1.81 74.76 +1.43
$(H_{\text{ub}}-w)L_{\text{KD}}$ 72.73 -0.60 68.90 -4.43
Leveraging this insight, adapting the focus of KD to valuable samples through entropy reweighting should fill the need for enhanced distillation. We conducted a preliminary experiment that compared the performance of reweighting with teacher ($H^{\mathcal{T}}$) and student ($H^{\mathcal{S}}$) entropy, along with their linear-inverted variants ($H_{\text{ub}}-w$), where high-entropy samples received lower weight. Here, $H_{\text{ub}}$ is the upper bound of entropy ($log100$ for CIFAR-100). As shown in Tab. 1, reweighting with either $H^{\mathcal{T}}$ or $H^{\mathcal{S}}$ significantly improved KD accuracy, while inverted reweighting led to decreased performance. This supports our hypothesis that focusing on samples with valuable knowledge enhances student learning, with $H^{\mathcal{T}}$ proving more effective due to the teacher’s more reliable assessment of sample value.
Exploring deeper, we observed that $H^{\mathcal{S}}$ exhibited increasing variability across training epochs for the top 10% $H^{\mathcal{T}}$ samples (Fig. 3). This suggests that while the teacher finds these samples valuable, potentially due to the inherent differences in architecture or capacity, the student’s assessment fluctuates during training. Some samples remain consistently challenging (high $H^{\mathcal{S}}$), while others become progressively simpler (low $H^{\mathcal{S}}$) over time. This misalignment reveals the limitation of reweighting solely with $H^{\mathcal{T}}$, as it remains constant throughout the epochs and thus fails to capture the student’s evolving learning process.
Figure 3: Box Plot of Student vs. Teacher Entropy for Top 10% High Entropy Samples Across Epochs. $H^{\mathcal{S}}$ shows significant variations for high $H^{\mathcal{T}}$ samples, highlighting the student’s different views from the teacher. EA-KD achieves improved mimicry of the teacher with more stable entropy and better alignment.
In light of the above analysis, we introduce Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that directs learning toward high-value samples by reweighting the distillation loss. Leveraging both teacher and student entropies, EA-KD emphasizes samples that encapsulate critical knowledge as identified by the teacher, while dynamically adapting to the student’s evolving needs. Intuitively, this method addresses a fundamental limitation in standard KD frameworks, which often overlook the unique learning value of each sample and tend to bias towards simpler knowledge ones. Furthermore, EA-KD can be seamlessly integrated into most KD frameworks, enhancing their performance with negligible computational cost (Fig. 1). Extensive experiments on image classification, object detection, and LLM distillation demonstrate its efficacy and versatility across diverse KD frameworks. Our main contributions are summarized as follows.
  • We propose EA-KD, a plug-and-play KD method that adaptively reweights the distillation loss to prioritize samples valuable for learning, enabling a more nuanced and efficient knowledge transfer process.
  • We show that high-entropy samples carry critical knowledge in KD, and propose an entropy-based reweighting factor that integrates both teacher and student entropy for a tailored and dynamic focus on these samples.
  • We show that EA-KD consistently enhances performance in both logit- and feature-based KD across CV and LLM tasks, achieving SOTA results with negligible cost.
2 Related Work
Logit and Feature Distillation. Logit distillation [9, 38, 11] aligns the softened output logits of the teacher and student, valued for its simplicity and broad applicability. On the other hand, feature distillation minimizes divergence in intermediate feature representations, offering enhanced learning but often with higher computational costs [26, 1, 19]. Both pathways have achieved SOTA performance across tasks and domains. However, most of them typically adopt a static distillation scheme, such as treating all samples uniformly. Adaptive distillation addresses this limitation by introducing more dynamic knowledge transfer processes [30, 16, 20, 22, 13].
Adaptive Distillation. These methods refine knowledge transfer by adapting to data characteristics, student capabilities, or training dynamics. For instance, LS [30] uses Z-score standardization on logits, allowing the student to focus on essential relationships in the teacher’s output. CTKD [16] and MKD [20] adjust the temperature parameter to meet student needs, with Instance-T [16] highlighting the effectiveness of sample-wise adjustments. RW-KD [22] improves distillation by using meta-learning to find optimal weights for each sample. However, these methods often increase complexity and computational costs [16, 20, 22] or are limited to a specific type of KD [30]. In contrast, EA-KD offers a straightforward, more interpretable approach that is versatile across diverse KD frameworks.
Entropy in KD. Previous work introduced an entropy-based metric to quantify knowledge in KD, measuring how much visual information is retained or discarded [2]. Inspired by this, we adopted the entropy of model output to identify valuable samples in KD. AKD [13], which also leverages entropy in KD, adjusts weights based on teacher entropy in multi-teacher distillation, placing greater emphasis on low-entropy predictions. On the contrary, our approach prioritizes high-entropy samples in single-teacher distillation. While low entropy may indicate reliable knowledge in multi-teacher schemes, we view lower entropy as simpler knowledge and is less beneficial for student learning in a single-teacher context, as demonstrated in our preliminary analysis and experiments in Sec. 1.
3 Methodology
3.1 Preliminaries
Information Theory. Entropy quantifies the uncertainty or information content of a random variable [28]. For a given sample $x_{n}$, the entropy $H_{n}$ is computed as follows:
$H_{n}=-\sum_{j=1}^{C}\sigma(z_{n,j})\log(\sigma(z_{n,j}))\,,$ (1)
where $\sigma(\cdot)$ denotes the softmax function, $z_{n,j}$ is the logit for class $j$ of sample $x_{n}$, and $C$ is the number of classes.
Knowledge Distillation. The goal of vanilla KD is to transfer the knowledge encapsulated in the teacher’s softened probability outputs to the student [9]. In classification tasks, the probabilities $p$ are softened using the temperature-scaled softmax function:
$p_{i}(T)=\sigma(z,T)_{i}=\frac{\exp(\frac{z_{i}}{T})}{\sum_{j=1}^{C}\exp(\frac{z_{j}}{T})}\,,$ (2)
where $p_{i}(T)$ denotes the softened probability for class $i$, and $\sigma(z_{i},T)$ is the temperature-scaled softmax function. The temperature $T$ controls the smoothness of the distribution, revealing the subtle inter-class relationships.
The core of KD is to minimize the Kullback-Leibler divergence (KLD) between the teacher’s and student’s softened probabilities, the KD loss is defined as:
$\begin{split}L_{\text{KD}}&=\text{KL}(p^{\mathcal{T}}(T)\|p^{\mathcal{S}}(T))\cdot T^{2}\\ &=\sum_{j=1}^{C}p^{\mathcal{T}}_{j}(T)\log\left(\frac{p^{\mathcal{T}}_{j}(T)}{p^{\mathcal{S}}_{j}(T)}\right)\cdot T^{2}\,,\end{split}$ (3)
where $p^{\mathcal{T}}$ and $p^{\mathcal{S}}$ denote the teacher’s and student’s softened outputs, respectively. For simplicity of theoretical analysis, we set $T=1$ in this section. $L_{\text{KD}}$ then simplifies to:
$\begin{split}L_{\text{KD}}=\sum_{j=1}^{C}p^{\mathcal{T}}_{j}\log\left(\frac{p^{\mathcal{T}}_{j}}{p^{\mathcal{S}}_{j}}\right).\end{split}$ (4)
Figure 4: Illustration of the Uniform Distillation Scheme in Standard KD (Left) and Entropy-based Adaptive Reweighting in EA-KD (Right). The uniform distillation scheme overlooks the learning value of individual samples, while EA-KD prioritizes valuable samples based on both teacher and student assessments.
3.2 EA-KD
Limitations in Standard KDs. Most logit- and feature-based KD methods typically apply a uniform distillation across all samples, overlooking their unique learning value. This oversight can cause the model to over-prioritize simpler samples at the expense of more valuable, high-entropy ones. Taking KLD—the main loss function for logit-based methods—as an example, consider a student initialized with a uniform distribution $p^{\mathcal{S}}$ where $p^{\mathcal{S}}_{i}=\frac{1}{C}\,\forall\,i$. For a low-entropy sample with a teacher output $p^{\mathcal{T}}_{\text{low}}$ where $p^{\mathcal{T}}_{\text{low},j}\approx 1$ and $p^{\mathcal{T}}_{\text{low},i}\approx 0\,(i\neq j)$, the KLD simplifies to:
$\begin{split}\text{KL}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})&\approx p^{\mathcal{T}}_{\text{low},j}\cdot\log\left(\frac{p^{\mathcal{T}}_{\text{low},j}}{p^{\mathcal{S}}_{j}}\right)\\ &=\log\left(\frac{1}{p^{\mathcal{S}}_{j}}\right)\\ &=\log(C).\end{split}$ (5)
For a high-entropy distribution $p^{\mathcal{T}}_{\text{high}}\approx\frac{1}{C}\,\forall\,i$, the KLD is:
$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}})&\approx\sum_{j=1}^{C}\frac{1}{C}\cdot\log\left(\frac{1/C}{1/C}\right)\\ &=0.\end{split}$ (6)
Thus, we obtain:
$\text{KLD}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})>\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}}).$ (7)
This inequality highlights KLD’s overemphasis on low-entropy samples, causing them to dominate the loss over high-entropy ones. Similarly, MSE loss in feature-based methods also biases learning toward high-magnitude activations, which typically correspond to low-entropy samples. Such imbalance shifts the learning focus toward simpler samples, hindering the transfer of knowledge crucial for learning, especially during early training stages (refer to the overshadowing effect in the left part of Fig. 4).
Entropy-Based Reweighting. As discussed in Sec. 1, entropy can serve as a measure of the sample learning value in KD. In practice, we soften the entropy with an alternative temperature $T^{\prime}$ to better capture the value, expressed as:
$H_{n}=-\sum_{j=1}^{C}p_{n,j}(T^{\prime})\log(p_{n,j}(T^{\prime}))\,,$ (8)
where $p_{n,j}(T^{\prime})$ is the temperature-scaled probability of class $j$ for sample $x_{n}$.
To effectively target and emphasize samples valuable for the student, EA-KD’s reweighting factor $w_{\text{EA}}$ is formulated using two components: a base term $w_{\text{base}}$ and an interaction term $w_{\text{interact}}$. The base term captures the inherent value of each sample based on the teacher’s output entropy, defined as:
$w_{\text{base},n}=H^{\mathcal{T}}_{n},\quad H^{\mathcal{T}}_{n}\in[0,H_{\text{ub}}]\,,$ (9)
where $H^{\mathcal{T}}_{n}$ is the entropy of the teacher’s prediction for sample $x_{n}$, and $H_{\text{ub}}=\log(C)$ denotes the upper bound of entropy for $C$ classes. The interaction term $w_{\text{interact}}$, on the other hand, captures the interplay between the teacher and student perspectives by taking the normalized product of their entropies:
$w_{\text{interact},n}=\frac{H^{\mathcal{T}}_{n}\cdot H^{\mathcal{S}}_{n}}{H_{\text{ub}}},\quad w_{\text{interact},n}\in[0,H_{\text{ub}}]\,.$ (10)
where $H^{\mathcal{S}}_{n}$ is the student’s entropy. Finally, the EA-KD reweighting factor $w_{\text{EA},n}$ is defined as the average of the base and interaction terms:
$w_{\text{EA},n}=\frac{w_{\text{base},n}+w_{\text{interact},n}}{2},\quad w_{\text{EA},n}\in[0,H_{\text{ub}}]\,.$ (11)
By integrating both the teacher’s evaluation and the student’s evolving understanding, $w_{\text{EA}}$ effectively prioritizes valuable samples while dynamically adjusting the focus throughout training.
Reformulation. To better illustrate the influence of the student’s perspective in $w_{\text{EA},n}$, we reformulate Eq. 11 as:
$\begin{split}w_{\text{EA},n}&=\frac{H^{\mathcal{T}}_{n}+\frac{H^{\mathcal{T}}_{n}\cdot H^{\mathcal{S}}_{n}}{H_{\text{ub}}}}{2}\\ &=\frac{1}{2}H^{\mathcal{T}}_{n}\left(1+\frac{H^{\mathcal{S}}_{n}}{H_{\text{ub}}}\right)\,.\end{split}$ (12)
This reformulation expresses $w_{\text{EA},n}$ as a product of the $H^{\mathcal{T}}_{n}$ and a scaling factor that depends on the $H^{\mathcal{S}}_{n}/H_{\text{ub}}$. In this way, the teacher’s assessment of sample value is adaptively adjusted based on the student’s perspective. Depending on the entropy values, $w_{\text{EA}}$ behaves as follows:
$w_{\text{EA},n}=$ $H_{\text{ub}}$ if $H^{\mathcal{T}}_{n}\rightarrow H_{\text{ub}}\land H^{\mathcal{S}}_{n}\rightarrow H_{\text{ub}}$ (13a)
$w_{\text{EA},n}=$ $\frac{H_{\text{ub}}}{2}$ if $H^{\mathcal{T}}_{n}\rightarrow H_{\text{ub}}\land H^{\mathcal{S}}_{n}\rightarrow 0$ (13b)
$w_{\text{EA},n}=$ 00 0 if $H^{\mathcal{T}}_{n}\rightarrow 0\quad\forall H^{\mathcal{S}}_{n}$ (13c)
$w_{\text{EA},n}=$ $w_{\text{EA},n}$ otherwise.
When the teacher considers a sample highly valuable for learning ($H^{\mathcal{T}}_{n}\rightarrow H_{\text{ub}}$), the scaling factor adjusts $w_{\text{EA},n}$ based on the student’s view. If the student aligns (Eq. (13a)), the weight is maximized to emphasize this sample. Conversely, if the student is confident (Eq. (13b)), the weight reduces to half for moderate focus. However, when the teacher considers the sample simple (Eq. (13c)), the weight remains low regardless of $H^{\mathcal{S}}_{n}$, as the teacher considers it lacks valuable knowledge.
Table 2: Results on CIFAR-100. Accuracy (%) of EA-methods vs. originals across teacher-student pairs, with relative improvements highlighted. Avg. $\Delta$ shows average improvement across methods and model pairs. Best results are bolded, and second-best are underlined.
Type Teacher ResNet32$\times$4 WRN-28-4 WRN-40-2 VGG13 VGG13 ResNet50 ResNet32$\times$4
79.42 78.60 75.61 74.64 74.64 79.34 79.42
Student ResNet8$\times$4 WRN-16-2 WRN-40-1 VGG8 MN-V2 MN-V2 SN-V2
72.50 73.26 71.98 70.36 64.60 64.60 71.82 Avg. $\Delta$
Logit KD [9] 73.33 75.04 73.54 72.98 67.37 67.35 74.45
EA-KD 75.46 +2.13 75.79 +0.75 74.38 +0.84 74.08 +1.10 69.17 +1.80 69.67 +2.32 75.91 +1.46 +1.48+
CTKD [16] 73.91 75.29 73.93 73.52 68.46 68.47 75.31
EA-CTKD 75.18 +1.27 75.72 +0.43 74.03 +0.10 73.79 +0.27 69.19 +0.73 69.38 +0.91 76.02 +0.71 +0.63+
DKD [38] 76.32 76.45 74.81 74.68 69.71 70.35 77.07
EA-DKD 76.80 +0.48 76.74 +0.29 74.98 +0.17 75.07 +0.39 70.39 +0.68 70.98 +0.63 77.72 +0.65 +0.47+
MLD [11] 77.08 76.83 75.35 75.18 70.57 71.04 78.44
EA-MLD 77.65 +0.57 77.47 +0.64 75.77 +0.42 75.28 +0.10 70.72 +0.15 71.43 +0.39 78.85 +0.41 +0.38+
MLD+LS [30] 78.28 77.20 75.56 75.22 70.94 71.19 78.76
EA-MLD+LS 78.38 +0.09 77.60 +0.39 75.78 +0.22 75.38 +0.16 70.67 -0.27 71.36 +0.17 79.13 +0.37 +0.16+
Feature ReviewKD [1] 75.63 76.39 74.45 74.45 70.37 69.89 77.78
EA-ReviewKD 76.10 +0.47 76.95 +0.56 75.43 +0.98 74.56 +0.11 70.55 +0.18 69.80 -0.09 78.22 +0.44 +0.38+
FCFD [19] 76.62 77.00 75.46 75.22 70.65 71.00 78.18
EA-FCFD 77.50 +0.88 77.15 +0.15 75.30 -0.16 75.36 +0.14 71.02 +0.37 71.97 +0.97 78.75 +0.56 +0.42+
Avg. $\Delta$ +0.84 +0.46 +0.37 +0.33 +0.52 +0.76 +0.66 +0.56+
Loss Integration. As shown in Fig. 4, the flexible nature of $w_{\text{EA}}$ allows it to be plug-and-play into most distillation frameworks by reweighting the contribution of each sample based on its learning value. For instance, when integrated with vanilla KD, the loss is defined as:
$\begin{split}L_{\text{EA-KD}}&=\sum_{n=1}^{N}w_{\text{EA},n}\cdot L_{\text{KD},n}\\ &=\sum_{n=1}^{N}w_{\text{EA},n}\cdot\text{KL}(p^{\mathcal{T}}_{n}\|p^{\mathcal{S}}_{n})\,,\end{split}$ (14)
where $L_{\text{KD},n}$ is the distillation loss for sample $x_{n}$, and $p^{\mathcal{T}}_{n}$ and $p^{\mathcal{S}}_{n}$ denote the teacher and student’s softened probability for $x_{n}$, respectively. Consequently, EA-KD introduces a more nuanced and dynamic knowledge transfer to standard KD frameworks, ensuring that valuable samples receive enhanced focus throughout training.
4 Experiments
4.1 Experimental Setup
Datasets. We used CIFAR-100 [12] (50k training, 10k validation images; 100 classes) and Tiny-ImageNet [14] (100k training, 5k validation images; 200 classes) to evaluate our method for image classification, and used MS-COCO [18] (118k training, 5k validation images; 80 classes) for object detection. Furthermore, for LLM distillation, we employed five instruction-following datasets following [7]: Dolly [4], SInst [34], Vicuna [3], S-NI [35], and UnNI [10].
Table 3: Results on Tiny-ImageNet. Top-1 (first row) and top-5 accuracy (second row) for both EA-methods and their original versions using ResNet32$\times$4 teacher and ResNet8$\times$4 student are reported.
Teacher Student KD EA-KD MLD EA-MLD MLD+LS EA-MLD+LS FCFD EA-FCFD Avg. $\Delta$
64.41 55.25 56.00 59.39 +3.39 61.91 62.65 +0.74 56.45 61.93 +5.48 60.12 60.51 +0.39 +2.50
85.07 79.62 79.64 82.40 +2.76 83.77 84.47 +0.70 81.15 84.35 +3.20 82.90 83.22 +0.32 +1.74
Table 4: Results on Tiny-ImageNet with Transformers. Accuracy (%) for EA-KD and KD using transformer-based teachers and a ResNet8$\times$4 student. Each result is averaged over three runs.
Teacher Student KD EA-KD
ViT-B 71.12 ResNet8$\times$4 55.25 54.70 57.53 +2.83
DeiT-B 85.55 56.30 58.36 +2.06
Swin-B 86.30 56.15 59.58 +3.43
Table 5: Results on MS-COCO. Performance of EA-KD and KD using Faster-RCNN with FPN. Metrics include AP (overall), AP50, and AP75.
R-101 & R-18 R-50 & MV2
AP AP50 AP75 AP AP50 AP75
Teacher 42.04 62.48 45.88 40.22 61.02 43.81
Student 33.26 53.61 35.26 29.47 48.87 30.90
KD 33.97 54.66 36.62 30.13 50.28 31.35
EA-KD 34.78 56.14 37.19 31.81 53.18 33.18
$\Delta$ +0.81 +1.48 +0.57 +1.68 +2.90 +1.83
KD Frameworks. We evaluated EA-KD across representative SOTA logit-based (KD [9], CTKD [16], DKD [38], MLD [11], MLD+LS [30]) and feature-based methods (ReviewKD [1], FCFD [19]), reweighting their distillation loss while preserving each framework’s original structure. For fair comparisons, no hyperparameter was tuned for the EA-variants, except for the KD weight of MLD+LS [30] was reduced from 9.0 to 4.0 to avoid over-penalizing.
Implementation Details. The EA-KD and other variants were evaluated on various CNNs (VGG [29], ResNet [8], WideResNet [36], MobileNet [27], ShuffleNet [23]), vision transformer teachers (ViT [6], DeiT [32], Swin [21]), and GPT2 [25] for LLM distillation. We used the training settings of [31] for vanilla KD and the respective settings of each KD framework for CIFAR-100 and Tiny-ImageNet, [1] for MS-COCO, and [37] for LLMs. We set $T^{\prime}=3$ based on our ablation study, except $T^{\prime}=2$ for transformer teachers. All results averaged over five runs, and analyses were conducted using a ResNet32$\times$4 teacher and a ResNet8$\times$4 student on CIFAR-100.
4.2 Results
CIFAR-100. Tab. 2 presents results of various EA-methods and their baselines on CIFAR-100. EA-KD consistently improves both logit- and feature-based KD frameworks across most teacher-student pairs, with an average gain of 0.56%. The logit-based EA-MLD+LS achieved SOTA results in most pairings, while EA-FCFD excelled in two heterogeneous pairs. Additionally, EA-MLD and EA-FCFD secured most second-best results, outperforming the previous SOTA, MLD+LS [30]. These findings highlight EA-KD’s broad applicability and effectiveness in enhancing knowledge transfer across diverse types of KD frameworks.
Tiny-ImageNet. Tab. 3 presents results on Tiny-ImageNet across SOTA frameworks enhanced with EA. For instance, EA-KD significantly improves top-1 accuracy by 3.39% and top-5 accuracy by 2.76% over standard KD, achieving performance close to FCFD with less than a 1% difference. These results highlight EA’s effectiveness on a larger and more class-diverse dataset. Further, Tab. 4 highlights EA-KD’s ability to outperform KD when distilling from transformer-based teachers to a CNN student. While standard KD stopped improving the student when Swin-B was used as the teacher, EA-KD continued to enhance student performance, mitigating the capacity gap by allowing better teachers to produce better students.
MS-COCO. Extending to object detection, Tab. 5 presents the performance of EA-KD and KD on the MS-COCO dataset. EA-KD consistently outperforms KD across AP metrics for both model pairings, with particularly notable gains in AP50. This highlights EA-KD’s ability to effectively guide the student with valuable knowledge in a more complex visual task, thereby enhancing its performance.
Table 6: LLM Distillation Results. Rouge-L [17] scores averaged over five seeds for KD, EA-KD, RKLD, and JSD on each dataset. SFT denotes a student fine-tuned on the dataset.
Dataset
Method Dolly SInst Vicuna S-NI UnNI Avg.
Teacher GPT2-XL 27.19 14.64 16.30 27.55 31.42 23.42
SFT GPT2-S 22.94 10.11 15.17 16.21 18.68 16.62
KD 24.54 10.43 15.66 17.24 20.28 17.63
RKLD 24.38 10.73 15.71 17.31 20.96 17.82
JSD 23.86 10.20 15.50 16.20 19.17 16.98
EA-KD 24.95 10.59 16.41 18.27 21.46 18.34
$\Delta$ to KD +0.41 +0.16 +0.75 +1.03 +1.18 +0.71
LLM Distillation. To explore the full potential of EA-KD, we extended our experiments to LLM distillation, applying the reweighting at the sequence level of the model outputs. As shown in Tab. 6, EA-KD consistently outperforms standard KD across datasets. Moreover, EA-KD also surpasses two recently emerging methods in this field, Reverse KLD (RKLD) and Jensen-Shannon divergence (JSD), which aim to address KLD’s limitation of forcing the student to cover all modes of the complex LLM teacher distribution [7]. In this scenario, EA-KD’s dynamic prioritization of the most critical knowledge performed effectively in such complexity. This highlights EA-KD’s versatility beyond visual tasks, showcasing its ability to target valuable knowledge across diverse domains.
4.3 Empirical Analysis
In this section, we conduct ablation studies to evaluate two key components of EA-KD: $T^{\prime}$ and the reweighting factor. We then compare EA-KD and vanilla KD from multiple perspectives, showing how EA-KD improves teacher-student alignment and class separability. Additionally, we discuss the synergy between EA and DKD, demonstrating enhanced robustness in both the loss surface and hyperparameter stability. Finally, we highlight EA-methods computational efficiency across diverse KD frameworks.
(a) KLD for Top and Bottom 10% Entropy Samples
(b) Raw KLD and Training Accuracy (%)
Figure 5: Comparison of KD and EA-KD During Training. EA-KD enhances focus on high entropy samples, leading to a more effective and efficient learning process, as reflected in the overall raw KLD and training accuracy.
Ablation Study. Two experiment were performed to evaluate the sensitivity of $T^{\prime}$ and the effect of each reweighting components in $w_{\text{EA}}$. (i) Tab. 7 shows that a $T^{\prime}$ of 3 consistently delivers optimal performance across model combinations, highlighting its robustness in reflecting sample value during distillation. (ii) As shown in Tab. 8, the single-sided factor $w_{\text{base}}$ notably boosts baseline performance by emphasizing valuable samples, while $w_{\text{interact}}$ integrates teacher-student dynamics but may overly rely on the student responses, leading to limited gains. By combining these strengths, $w_{\text{EA}}$ achieves superior improvement by prioritizing valuable samples and adapting to the student’s learning dynamics, enabling a precise and tailored focus.
Table 7: Impact of $T^{\prime}$. Setting $T^{\prime}$ to 3 yields the best results across architectures, demonstrating its robustness and consistency.
Teacher Student $T^{\prime}$
2 3 4
ResNet32$\times$4 ResNet8$\times$4 73.90 75.46 75.22
WRN-28-4 WRN-16-2 74.89 75.79 75.62
ResNet32$\times$4 SN-V2 75.51 75.91 75.72
Table 8: Impact of Reweighting Factors. The proposed $w_{\text{EA}}$ outperforms the single-sided $w_{\text{base}}$, while $w_{\text{interact}}$, lacking teacher base guidance, shows limited improvement.
Method Acc. Reweighting Factor
$w_{\text{base}}$ $w_{\text{interact}}$ $w_{\text{EA}}$
KD [9] 73.33 $\uparrow$ 75.14 $\uparrow$ $\uparrow$ 74.76 $\uparrow$ $\uparrow$ 75.27 $\uparrow$
MLD [11] 77.08 $\uparrow$ 77.47 $\uparrow$ $\uparrow$ 77.45 $\uparrow$ $\uparrow$ 77.65 $\uparrow$
MLD+LS [30] 78.28 $\uparrow$ 78.30 $\uparrow$ $\uparrow$ 78.20 $\downarrow$ $\uparrow$ 78.38 $\uparrow$
FCFD [19] 76.62 $\uparrow$ 77.37 $\uparrow$ $\uparrow$ 77.42 $\uparrow$ $\uparrow$ 77.44 $\uparrow$
Refined Knowledge Focus and Alignment. Fig. 5(a) illustrates the KLD for the top and bottom 10% entropy samples throughout training for both KD and EA-KD, with Fig. 5(b) depicting the overall raw KLD and training accuracy. Notably, vanilla KD’s uniform distillation causes low-entropy samples to dominate the distillation loss, failing to transfer the more critical knowledge from higher-entropy samples, as discussed in Sec. 3. Fig. 2(a) further illustrates this, where KD shows increasingly larger accuracy gaps with the teacher on higher entropy samples. In contrast, EA-KD refined the student’s focus on high-entropy samples and efficiently reduced the raw KLD (Fig. 5), achieving improved mimicry and generalization with more consistent alignment to the teacher across entropy segments (Fig. 2(a)).
Student’s Unique Perspective. Fig. 6 compares the t-SNE visualizations [33] of KD and EA-KD with the high $H^{\mathcal{S}}$ and $w_{\text{EA}}$ samples highlighted. Similar to Fig. 2(b), high $H^{\mathcal{S}}$ samples also lie near decision boundaries in the KD-student. This suggests that, despite having a different view from the teacher (Fig. 3), the student’s evolving capacity enables $H^{\mathcal{S}}$ to capture the samples crucial for its own learning progress [^3] See Sec. A.1 for further t-SNE visualizations showing the student’s evolving sample focus over training epochs.. In EA-KD, integrating both $H^{\mathcal{T}}$ and $H^{\mathcal{S}}$ in $w_{\text{EA}}$ allows us to harness the complementary strengths of the teacher’s informed assessment and the student’s dynamic learning needs. Consequently, EA-KD achieves superior class separability as indicated by a higher Calinski-Harabasz (CH) index.
Consistency in Entropy Levels. As shown in Fig. 3, EA-KD mitigates the increasing entropy variation across epochs observed in KD, enabling the student to maintain more stable $H^{\mathcal{S}}$ and align more closely to $H^{\mathcal{T}}$. This stability creates a feedback loop in EA-KD, where a steady $H^{\mathcal{S}}$ leads to a stable $w_{\text{EA}}$, promoting a more focused learning process. As a result, the student more effectively mimics the teacher’s response and achieves improved performance.
Figure 6: t-SNE of Students. Top 10% high $H^{\mathcal{S}}$ samples for KD (left) and $w_{\text{EA}}$ samples for EA-KD (right) are highlighted with triangles. High $H^{\mathcal{S}}$ samples cluster near decision boundaries, underscoring their learning value for the student. EA-KD achieves better class separability as indicated by a higher CH index.
Figure 7: Loss Surface and Differences in Area@1.6 for DKD and EA-DKD Students Across Epochs. The mean and variance for the surface, along with the contour areas at level 1.6 (Area@1.6), are provided for each subplot. The line plot (lower right) tracks the differences in Area@1.6 between the teacher and students over epochs. EA-DKD consistently shows larger contours with less fluctuation, signifying smoother learning surfaces and a more robust generalization process compared to DKD.
Figure 8: Comparison of DKD and EA-DKD Performance Across Varying $\beta$ Values. EA-DKD consistently outperforms DKD and significantly reduces performance variance over $\beta$ (0.10 vs. 0.31), highlighting EA-DKD’s enhanced robustness.
Loss Landscape and Synergy in EA-DKD. Visualizing the loss landscape [15] offers insights into the student’s robustness against noise and generalization. As illustrated in Fig. 7, DKD shows narrow and fluctuating contours with significantly higher variance across epochs (e.g. 366.26 at epoch 240), indicating instability in the loss surface and weaker generalization. In contrast, EA-DKD consistently produces a smoother surface with more stable contours (e.g. variance of 178.20 at epoch 240) and broader low-loss regions (Area@1.6) compared to DKD, suggesting improved generalizability. In addition, the Area@1.6 comparison (Fig. 7, lower right) underscores EA-DKD’s closer alignment with the teacher’s surface throughout training.
The enhanced generalization in EA-DKD also extends to the robustness of its hyperparameter. DKD was introduced to decouple target class KD (TCKD) and non-target class KD (NCKD) in vanilla KD, balancing them with hyperparameters $\alpha$ and $\beta$ [38]. While $\alpha$ remains stable around 1.0, adjusting $\beta$ from 1.0 to 10.0 leads to fluctuating performance [38], with a variance of 0.31 (see Fig. 8). Notably, EA-DKD effectively reduces this fluctuation to a variance of 0.10 and consistently improves the performance. Overall, the enhanced generalizability and robustness may be attributed to the synergy between DKD and EA in handling different aspects of knowledge transfer: DKD prevents NCKD from being overshadowed by TCKD at the class level, while EA ensures valuable samples—often rich in NCKD—are not dominated by simpler ones at the sample level. Additionally, the dynamic focus in EA compensates for the static nature of $\beta$. Together, EA-DKD facilitates a more balanced and efficient knowledge transfer process, both class-wise and sample-wise.
Computational Efficiency. Fig. 1 illustrates the efficiency of EA-methods, showing significant performance improvements on various SOTA KD frameworks with negligible cost. This remarkable computational efficiency, combined with its streamlined integration into both logit- and feature-based KD frameworks, suggests the potential of our method as a versatile and accessible enhancement for KD.
5 Conclusion
In this paper, we examined existing KD methods from a novel perspective, revealing the limitation of their inherent uniform distillation scheme, which often hinders the transfer of valuable, high-entropy samples that contain critical knowledge. To overcome this, we introduced EA-KD, a plug-and-play KD approach that dynamically reweights the distillation loss, directing the learning process toward valuable samples. EA-KD consistently improves SOTA KD methods across image classification, object detection, and LLM distillation, all with negligible cost. We believe EA-KD showcases a great paradigm of the meticulous handling of knowledge transfer, offering sample-wise adaptability while integrating the student’s evolving learning dynamics into the distillation process.
Supplementary Material
Appendix A Additional Analysis and Visualizations
This document provides additional analysis to complement the main paper. It includes a detailed examination of high student entropy samples, loss landscape comparisons for KD and EA-KD, and t-SNE visualizations for various KD frameworks and their EA-enhanced variants.
A.1 Analysis of High Student Entropy Samples
Complementary to the teacher entropy analysis in Sec. 1, this section emphasizes the critical role of high student entropy samples in the adaptive reweighting process of EA-KD. As shown in Fig. 1(a), high student entropy samples correlate with larger teacher-student accuracy gaps similar to the teacher entropy segments (Fig. 2(a)). Additionally, Fig. 1(b) presents the students’ t-SNE visualizations across training epochs for KD and EA-KD, with high-entropy samples ($H^{\mathcal{S}}$) and reweighted samples ($w_{\text{EA}}$) highlighted, respectively. Notably, these high-entropy samples also cluster near decision boundaries similar to the teacher’s t-SNE (Fig. 2(b)). Unlike the teacher, however, the student’s top-entropy samples shift dynamically over training. For instance, in KD, the yellow class at epoch 40 is located near the center of most clusters and contains numerous high-entropy samples. By epoch 120, high-entropy samples are more prevalent in the green and purple classes as they shift closer to the center. With this dynamic integrated into $w_{\text{EA}}$, EA-KD continuously adapts to the student’s evolving learning needs throughout training, complementing the static nature of $w_{\text{base}}$ and fostering a tailored knowledge transfer.
(a) Accuracy vs. Student Entropy Segments.
(b) t-SNE of Students Across Training Epochs.
Figure A1: High Student Entropy Analysis in KD and EA-KD. (a) Higher student entropy samples correlate with larger accuracy gaps in KD. EA-KD demonstrates closer alignment with the teacher, particularly in high-entropy regions. (b) High $H^{\mathcal{S}}$ samples consistently cluster near decision boundaries, reflecting the student’s real-time learning needs. EA-KD adapts to this dynamic, achieving enhanced class separability over epochs.
A.2 Loss Landscape Analysis for KD and EA-KD
We further analyze the robustness of KD and EA-KD by examining their loss landscapes across training epochs. As shown in Fig. A2, EA-KD achieves a consistently smoother loss surface compared to KD (e.g. $\sigma^{2}$ of 217.68 vs. 305.21 at epoch 240) and larger Area@1.6 values. Notably, KD’s Area@1.6 metric remains 0 at epochs 40 and 120, indicating the absence of low-loss regions. In contrast, EA-KD achieves values of 0.01 and 0.05, respectively, underscoring its improved learning efficiency by enabling the student to assimilate the key knowledge. These results highlight EA-KD’s ability to ensure a more efficient and effective optimization, facilitating enhanced generalizability compared to standard KD.
A.3 t-SNE of KD Frameworks and EA-methods
The t-SNE visualizations of students from various KD frameworks and their EA-method variants, alongside the teacher’s representation, are shown in Fig. A3. EA-methods consistently exhibit more distinct and well-defined class clusters compared to their respective baselines, as evidenced by higher CH indices. Remarkably, the SOTA EA-MLD+LS achieves the highest CH index of 814.96, indicating its superior performance and closer alignment with the teacher. Furthermore, although EA-DKD ranks fourth in performance (Tab. 2), it demonstrates the second-best class separability, underscoring its enhanced robustness as discussed in Sec. 4.3. These findings highlight EA-KD’s versatility in enhancing class separability and improving knowledge transfer across diverse KD frameworks.
Figure A2: Loss Surface and Differences in Area@1.6 for KD and EA-KD Students Across Epochs. Contour plots illustrate the loss landscapes of KD (first row) and EA-KD (second row) across training epochs. The line plot (lower right) tracks the differences in Area@1.6 between the teacher and students over epochs. EA-KD exhibits a more stable and robust loss surface, with greater Area@1.6 earlier in training, signifying a more efficient learning process.
Figure A3: t-SNE Visualizations of EA-methods vs. Baselines. The students from various KD frameworks (first and third row) and their EA-enhanced counterparts (second and fourth row), along with the teacher (lower right), are shown. EA-methods consistently achieve higher CH indices, indicating better class separability.

### v3

@@
 In this paper, we propose Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that prioritizes learning from valuable samples.
-EA-KD quantifies each sample’s learning value by strategically combining the entropy of the teacher and student output, then dynamically reweights the distillation loss to place greater emphasis on high-value samples.
+EA-KD quantifies each sample’s learning value by strategically combining the entropy of the teacher and student output, then dynamically reweights the distillation loss to place greater emphasis on high-entropy samples.
 Extensive experiments across diverse KD frameworks and tasks–including image classification, object detection, and large language model (LLM) distillation–demonstrate that EA-KD consistently enhances performance, achieving state-of-the-art results with negligible computational cost.
-Our code will be publicly available.
-11footnotetext: Equal contribution22footnotetext: Corresponding author: camhero@gmail.com
+Our code is available at https://github.com/cpsu00/EA-KD.
+11footnotetext: Equal contribution.22footnotetext: Bin Pu and Shin-Jye Lee are corresponding authors (eebinpu@ust.hk, camhero@gmail.com).
 1 Introduction
-The increasing size of state-of-the-art (SOTA) deep learning models in computer vision (CV) [32, 21, 6], and large language models (LLMs) [5, 25] pose challenges for deployment in resource-constrained settings.
-Knowledge Distillation (KD) [9] offers a solution by training a smaller “student” model to mimic a larger “teacher” model, using both ground-truth labels and the teacher’s “knowledge” (e.g.
-softened logits or feature representations) to achieve similar performance in a compact form.
-Figure 1: Training Time per Batch (ms) vs. Testing Accuracy (%) on CIFAR-100. Highlighting EA methods’ enhanced performance with minimal additional training cost.
-SOTA KD methods have explored diverse knowledge forms and structural modifications to refine knowledge transfer [9, 26, 1, 24, 38, 11, 19].
-However, most methods distill uniformly across all samples, operating under the assumption that each sample has equal importance and overlooking their varying learning needs.
+The growing size of state-of-the-art (SOTA) deep learning models poses challenges for deployment in resource-constrained settings.
+Knowledge Distillation (KD) [11] offers a solution by training a smaller “student” model to mimic a larger “teacher” model, using both ground-truth and the teacher’s “knowledge” (e.g.
+logits or feature representations) to achieve similar performance in a compact form.
+KD has been widely applied across domains, including computer vision (CV) [40, 1, 31], large language models (LLMs) [35, 13, 9, 15], and medical imaging [21, 22, 23, 8].
+Figure 1: Training Time per Batch (ms) vs. Accuracy (%) of EA-Methods and Baselines on CIFAR-100.
+Advanced KD methods have explored diverse knowledge forms and structural modifications to refine knowledge transfer [11, 33, 2, 30, 47, 14, 26].
+However, most methods distill uniformly across all samples, operating under the assumption that each sample has equal importance and overlooking their varying learning value.
 Recent advancements, such as Instance-T [20], have shown that assigning unique temperatures to each sample outperforms the uniform temperature method Global-T.
@@
 can further optimize the distillation process.
-This mimics how human students learn better when key points are highlighted by teachers, and thus, this targeted approach allows the student model to achieve better comprehension by focusing on more informative samples.
-As we will demonstrate, the uniform distillation in most KD methods also tends to overshadow these important samples, limiting the efficiency of knowledge transfer.
+This mirrors how human students learn better when key points are highlighted by teachers.
+Such a strategy enables the student model to focus on more informative samples, leading to improved performance.
+However, as we will show, the uniform distillation scheme in most KD methods often overlooks these critical samples, thereby limiting the efficiency of knowledge transfer.
 The question then arises: How can we identify the most valuable samples for learning?
 Entropy, the core concept in information theory that quantifies the uncertainty or information of a random variable [36], may be well-suited for this role.
-While a prior multi-teacher KD study utilized inverse entropy to measure the reliability of teachers’ output [13], this method does not apply to the more prevalent single-teacher setting.
-Moreover, relying solely on the teacher’s entropy restricts the adaptability to the evolving learning dynamics of the student, leaving the full potential of entropy in KD under-explored.
+Prior methods have incorporated entropy in KD to adjust weighting [17] or refine logit predictions [49]; however, these approaches are typically restricted to specific scenarios (e.g., multi-teacher or logit-based KD) and primarily rely on teacher entropy, leaving the broader potential of entropy in KD underexplored.
 Therefore, we propose leveraging entropy to quantify the learning value of each sample in KD, as high-entropy outputs shall correspond to greater information contents that are crucial for learning.
 Our preliminary analysis reveals that higher entropy samples [^2] Results for high teacher entropy samples are shown here; high student entropy plots are in the Sec.
-A.1.
+B.1.
 (i) correlate with larger teacher-student accuracy gaps (Fig.
@@
 2(b)), suggesting these informative samples not only offer valuable learning opportunities but are also pivotal in defining decision boundaries.
-Hence, entropy can serve as a reliable metric for identifying valuable samples that are crucial for learning in KD.
+Thus, entropy can serve as a reliable metric for identifying the most valuable samples in KD.
 (a) Accuracy vs. Entropy Segments.
@@
 (a) Higher teacher entropy correlates with larger accuracy gaps in KD, while EA-KD maintains closer alignment.
-(b) The top 10% of high entropy samples (denoted by triangles) cluster near decision boundaries, representing critical knowledge essential for classification.
+(b) The top 10% high entropy samples (denoted by triangles) cluster near decision boundaries, representing critical knowledge essential for classification.
 Table 1: Entropy-based Reweighting on KD.
@@
 $(H_{\text{ub}}-w)L_{\text{KD}}$ 72.73 -0.60 68.90 -4.43
-Leveraging this insight, adapting the focus of KD to valuable samples through entropy reweighting should fill the need for enhanced distillation.
-We conducted a preliminary experiment that compared the performance of reweighting with teacher ($H^{\mathcal{T}}$) and student ($H^{\mathcal{S}}$) entropy, along with their linear-inverted variants ($H_{\text{ub}}-w$), where high-entropy samples received lower weight.
-Here, $H_{\text{ub}}$ is the upper bound of entropy ($log100$ for CIFAR-100).
+Leveraging this insight, adapting the focus of KD to valuable samples should fill the need for enhanced distillation.
+To validate this, we conducted a preliminary experiment that compared the performance of reweighting with teacher ($H^{\mathcal{T}}$) and student ($H^{\mathcal{S}}$) entropy, along with their linear-inverted variants ($H_{\text{ub}}-w$), where high-entropy samples received lower weight.
+Here, $H_{\text{ub}}$ denotes the upper bound of entropy (i.e., $\log 100$ for CIFAR-100).
 As shown in Tab.
@@
 This misalignment reveals the limitation of reweighting solely with $H^{\mathcal{T}}$, as it remains constant throughout the epochs and thus fails to capture the student’s evolving learning process.
-Figure 3: Box Plot of Student vs.
-Teacher Entropy for Top 10% High Entropy Samples Across Epochs.
-$H^{\mathcal{S}}$ shows significant variations for high $H^{\mathcal{T}}$ samples, highlighting the student’s different views from the teacher.
-EA-KD achieves improved mimicry of the teacher with more stable entropy and better alignment.
-In light of the above analysis, we introduce Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that directs learning toward high-value samples by reweighting the distillation loss.
+Figure 3: Box Plot of $H^{\mathcal{S}}$ vs.
+$H^{\mathcal{T}}$ for Top 10% High $H^{\mathcal{T}}$ Samples.
+$H^{\mathcal{S}}$ shows notable variation, reflecting the student’s learning dynamics and differing perspective from the teacher.
+EA-KD enhances mimicry with more stable and aligned entropy.
+In light of the above analysis, we introduce Entropy-based Adaptive Knowledge Distillation (EA-KD), a simple yet effective plug-and-play KD method that enhances distillation by reweighting the loss toward high-value samples.
 Leveraging both teacher and student entropies, EA-KD emphasizes samples that encapsulate critical knowledge as identified by the teacher, while dynamically adapting to the student’s evolving needs.
-Intuitively, this method addresses a fundamental limitation in standard KD frameworks, which often overlook the unique learning value of each sample and tend to bias towards simpler knowledge ones.
+This method directly addresses a fundamental limitation in standard KD frameworks, which often overlook the unique learning value of each sample and tend to bias towards simpler knowledge.
 Furthermore, EA-KD can be seamlessly integrated into most KD frameworks, enhancing their performance with negligible computational cost (Fig.
@@
 Extensive experiments on image classification, object detection, and LLM distillation demonstrate its efficacy and versatility across diverse KD frameworks.
-Our main contributions are summarized as follows.
-  • We propose EA-KD, a plug-and-play KD method that adaptively reweights the distillation loss to prioritize samples valuable for learning, enabling a more nuanced and efficient knowledge transfer process.
-  • We show that high-entropy samples carry critical knowledge in KD, and propose an entropy-based reweighting factor that integrates both teacher and student entropy for a tailored and dynamic focus on these samples.
-  • We show that EA-KD consistently enhances performance in both logit- and feature-based KD across CV and LLM tasks, achieving SOTA results with negligible cost.
+Our main contributions are summarized as follows:
+  • We reveal that high-entropy samples carry critical knowledge in KD and propose an entropy-based reweighting factor that integrates both teacher and student entropy to provide a dynamic and tailored learning focus.
+  • We introduce EA-KD, a plug-and-play KD method that adaptively reweights the distillation loss to prioritize valuable samples, enabling more effective and efficient knowledge transfer.
+  • We demonstrate that EA-KD consistently improves performance across logit- and feature-based KD methods, achieving SOTA results on both CV and LLM tasks with minimal computational overhead.
 2 Related Work
@@
 However, most of them typically adopt a static distillation scheme, such as treating all samples uniformly.
-Adaptive distillation addresses this limitation by introducing more dynamic knowledge transfer processes [30, 16, 20, 22, 13].
+Adaptive distillation addresses this limitation by introducing more dynamic knowledge transfer processes [28, 46, 20, 38, 17, 49, 48]
 Adaptive Distillation.
-These methods refine knowledge transfer by adapting to data characteristics, student capabilities, or training dynamics.
-For instance, LS [30] uses Z-score standardization on logits, allowing the student to focus on essential relationships in the teacher’s output.
-CTKD [16] and MKD [20] adjust the temperature parameter to meet student needs, with Instance-T [16] highlighting the effectiveness of sample-wise adjustments.
-RW-KD [22] improves distillation by using meta-learning to find optimal weights for each sample.
-However, these methods often increase complexity and computational costs [16, 20, 22] or are limited to a specific type of KD [30].
-In contrast, EA-KD offers a straightforward, more interpretable approach that is versatile across diverse KD frameworks.
+These methods improve knowledge transfer by dynamically adjusting knowledge at different levels.
+For sample-level, RW-KD [28] employs meta-learning to optimize weights for each sample, which introduces high computational overhead.
+PAD [46] showed that traditional hard-mining weighting is unsuitable for KD and instead prioritizes samples with low uncertainty and small teacher-student gaps.
+This shares similarities with EA-KD, as we will show that high-entropy samples often exhibit lower KLD.
+However, PAD relies on an auxiliary estimation branch, whereas EA-KD offers a more interpretable and efficient approach by directly utilizing entropy.
+For logit-level, CTKD [20] and LS [38] dynamically adjust the temperature parameter $T$ to refine knowledge transfer, but remain limited to logit-based KD.
+Instead, EA-KD’s sample-wise reweighting ensures broader compatibility across KD frameworks.
+Importantly, EA-KD and these methods serve distinct yet complementary roles, combining them could further improve performance, as we will show.
 Entropy in KD.
-Previous work introduced an entropy-based metric to quantify knowledge in KD, measuring how much visual information is retained or discarded [2].
-Inspired by this, we adopted the entropy of model output to identify valuable samples in KD.
-AKD [13], which also leverages entropy in KD, adjusts weights based on teacher entropy in multi-teacher distillation, placing greater emphasis on low-entropy predictions.
-On the contrary, our approach prioritizes high-entropy samples in single-teacher distillation.
-While low entropy may indicate reliable knowledge in multi-teacher schemes, we view lower entropy as simpler knowledge and is less beneficial for student learning in a single-teacher context, as demonstrated in our preliminary analysis and experiments in Sec.
-1.
+Cheng et al.
+introduced an entropy-based metric to quantify knowledge retention in KD [3].
+Inspired by this, we utilize entropy to identify valuable samples in KD.
+AKD [17] assign higher weights to low-entropy teacher predictions in multi-teacher settings.
+However, such weighting can degrade performance in the more common single-teacher settings (Tab.
+1).
+DynamicKD [49] refines logit-level knowledge through entropy correction similar to CTKD and LS, and is also constrained to logit-based KD.
+TTM [48] removes the student temperature, revealing an inherent Rényi entropy regularization, while WTTM further emphasizes the uncertain samples.
+In contrast, EA-KD actively leverages both teacher and student entropy for dynamic weighting, yielding stronger performance [^3] See Sec.
+A.1 for a more detailed comparison..
 3 Methodology
@@
 Information Theory. Entropy quantifies the uncertainty or information content of a random variable [36]. For a given sample $x_{n}$, the entropy $H_{n}$ is computed as follows:
-$H_{n}=-\sum_{j=1}^{C}\sigma(z_{n,j})\log(\sigma(z_{n,j}))\,,$ (1)
-where $\sigma(\cdot)$ denotes the softmax function, $z_{n,j}$ is the logit for class $j$ of sample $x_{n}$, and $C$ is the number of classes.
+$H_{n}=-\sum_{i=1}^{C}\sigma(z_{n,i})\log(\sigma(z_{n,i}))\,$, (1)
+where $\sigma(\cdot)$ denotes the softmax function, $z_{n,i}$ is the logit for class $i$ of sample $x_{n}$, and $C$ is the number of classes.
 Knowledge Distillation.
@@
 In classification tasks, the probabilities $p$ are softened using the temperature-scaled softmax function:
-$p_{i}(T)=\sigma(z,T)_{i}=\frac{\exp(\frac{z_{i}}{T})}{\sum_{j=1}^{C}\exp(\frac{z_{j}}{T})}\,,$ (2)
+$p_{i}(T)=\sigma(z,T)_{i}=\frac{\exp(\frac{z_{i}}{T})}{\sum_{k=1}^{C}\exp(\frac{z_{k}}{T})}\,$, (2)
 where $p_{i}(T)$ denotes the softened probability for class $i$, and $\sigma(z_{i},T)$ is the temperature-scaled softmax function.
@@
 The core of KD is to minimize the Kullback-Leibler divergence (KLD) between the teacher’s and student’s softened probabilities, the KD loss is defined as:
-$\begin{split}L_{\text{KD}}&=\text{KL}(p^{\mathcal{T}}(T)\|p^{\mathcal{S}}(T))\cdot T^{2}\\ &=\sum_{j=1}^{C}p^{\mathcal{T}}_{j}(T)\log\left(\frac{p^{\mathcal{T}}_{j}(T)}{p^{\mathcal{S}}_{j}(T)}\right)\cdot T^{2}\,,\end{split}$ (3)
+[…]
 where $p^{\mathcal{T}}$ and $p^{\mathcal{S}}$ denote the teacher’s and student’s softened outputs, respectively.
@@
 $L_{\text{KD}}$ then simplifies to:
-$\begin{split}L_{\text{KD}}=\sum_{j=1}^{C}p^{\mathcal{T}}_{j}\log\left(\frac{p^{\mathcal{T}}_{j}}{p^{\mathcal{S}}_{j}}\right).\end{split}$ (4)
-Figure 4: Illustration of the Uniform Distillation Scheme in Standard KD (Left) and Entropy-based Adaptive Reweighting in EA-KD (Right).
-The uniform distillation scheme overlooks the learning value of individual samples, while EA-KD prioritizes valuable samples based on both teacher and student assessments.
+$\begin{split}L_{\text{KD}}=\sum_{i=1}^{C}p^{\mathcal{T}}_{i}\log\left(\frac{p^{\mathcal{T}}_{i}}{p^{\mathcal{S}}_{i}}\right).\end{split}$ (4)
+Figure 4: Illustration of the Uniform Distillation Scheme in Standard KD (Left) and Entropy-based Reweighting in EA-KD (Right).
+Standard KD not only overlook the varying learning value of individual samples but also bias toward low-entropy (easier) ones.
+In contrast, EA-KD effectively guides learning toward valuable samples based on both teacher and student assessments.
 3.2 EA-KD
 Limitations in Standard KDs.
-Most logit- and feature-based KD methods typically apply a uniform distillation across all samples, overlooking their unique learning value.
+Most logit- and feature-based KD methods [11, 47, 14, 2, 26, 33] treat all samples uniformly, overlooking their unique learning value.
 This oversight can cause the model to over-prioritize simpler samples at the expense of more valuable, high-entropy ones.
 Taking KLD—the main loss function for logit-based methods—as an example, consider a student initialized with a uniform distribution $p^{\mathcal{S}}$ where $p^{\mathcal{S}}_{i}=\frac{1}{C}\,\forall\,i$.
-For a low-entropy sample with a teacher output $p^{\mathcal{T}}_{\text{low}}$ where $p^{\mathcal{T}}_{\text{low},j}\approx 1$ and $p^{\mathcal{T}}_{\text{low},i}\approx 0\,(i\neq j)$, the KLD simplifies to:
-$\begin{split}\text{KL}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})&\approx p^{\mathcal{T}}_{\text{low},j}\cdot\log\left(\frac{p^{\mathcal{T}}_{\text{low},j}}{p^{\mathcal{S}}_{j}}\right)\\ &=\log\left(\frac{1}{p^{\mathcal{S}}_{j}}\right)\\ &=\log(C).\end{split}$ (5)
+For a low-entropy sample with a teacher output $p^{\mathcal{T}}_{\text{low}}$ where $p^{\mathcal{T}}_{\text{low},j}\approx 1$ and $p^{\mathcal{T}}_{\text{low},i}\approx 0\,(i\neq j)$, the KLD becomes:
+$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})&\approx p^{\mathcal{T}}_{\text{low},j}\cdot\log\left(\frac{p^{\mathcal{T}}_{\text{low},j}}{p^{\mathcal{S}}_{j}}\right)\\ &=\log\left(\frac{1}{p^{\mathcal{S}}_{j}}\right)\\ &=\log(C).\end{split}$ (5)
 For a high-entropy distribution $p^{\mathcal{T}}_{\text{high}}\approx\frac{1}{C}\,\forall\,i$, the KLD is:
-$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}})&\approx\sum_{j=1}^{C}\frac{1}{C}\cdot\log\left(\frac{1/C}{1/C}\right)\\ &=0.\end{split}$ (6)
+$\begin{split}\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}})&\approx\sum_{i=1}^{C}\frac{1}{C}\cdot\log\left(\frac{1/C}{1/C}\right)\\ &=0.\end{split}$ (6)
 Thus, we obtain:
 $\text{KLD}(p^{\mathcal{T}}_{\text{low}}\|p^{\mathcal{S}})>\text{KLD}(p^{\mathcal{T}}_{\text{high}}\|p^{\mathcal{S}}).$ (7)
-This inequality highlights KLD’s overemphasis on low-entropy samples, causing them to dominate the loss over high-entropy ones.
-Similarly, MSE loss in feature-based methods also biases learning toward high-magnitude activations, which typically correspond to low-entropy samples.
+This inequality highlights KLD’s inherent bias toward low-entropy samples, which dominate the loss over the valuable, high-entropy ones.
+Similarly, the MSE loss in feature-based KD also biases learning toward high-magnitude activations, which typically correspond to low-entropy samples.
 Such imbalance shifts the learning focus toward simpler samples, hindering the transfer of knowledge crucial for learning, especially during early training stages (refer to the overshadowing effect in the left part of Fig.
@@
 1, entropy can serve as a measure of the sample learning value in KD.
-In practice, we soften the entropy with an alternative temperature $T^{\prime}$ to better capture the value, expressed as:
-$H_{n}=-\sum_{j=1}^{C}p_{n,j}(T^{\prime})\log(p_{n,j}(T^{\prime}))\,,$ (8)
-where $p_{n,j}(T^{\prime})$ is the temperature-scaled probability of class $j$ for sample $x_{n}$.
-To effectively target and emphasize samples valuable for the student, EA-KD’s reweighting factor $w_{\text{EA}}$ is formulated using two components: a base term $w_{\text{base}}$ and an interaction term $w_{\text{interact}}$.
+In practice, we soften the entropy with an alternative temperature $T^{\prime}$ to better reflect sample value:
+$H_{n}=-\sum_{i=1}^{C}p_{n,i}(T^{\prime})\log(p_{n,i}(T^{\prime}))\,$, (8)
+where $p_{n,i}(T^{\prime})$ is the temperature-scaled probability of class $i$ for sample $x_{n}$.
+To dynamically emphasize valuable samples, EA-KD’s reweighting factor $w_{\text{EA}}$ is formulated using two components: a base term $w_{\text{base}}$ and an interaction term $w_{\text{interact}}$.
 The base term captures the inherent value of each sample based on the teacher’s output entropy, defined as:
 $w_{\text{base},n}=H^{\mathcal{T}}_{n},\quad H^{\mathcal{T}}_{n}\in[0,H_{\text{ub}}]\,$, (9)
-where $H^{\mathcal{T}}_{n}$ is the entropy of the teacher’s prediction for sample $x_{n}$, and $H_{\text{ub}}=\log(C)$ denotes the upper bound of entropy for $C$ classes.
+where $H^{\mathcal{T}}_{n}$ denotes the entropy of the teacher’s prediction for sample $x_{n}$, and $H_{\text{ub}}=\log(C)$ is the upper bound of entropy for $C$ classes.
 The interaction term $w_{\text{interact}}$, on the other hand, captures the interplay between the teacher and student perspectives by taking the normalized product of their entropies:
@@
 $\begin{split}w_{\text{EA},n}&=\frac{H^{\mathcal{T}}_{n}+\frac{H^{\mathcal{T}}_{n}\cdot H^{\mathcal{S}}_{n}}{H_{\text{ub}}}}{2}\\ &=\frac{1}{2}H^{\mathcal{T}}_{n}\left(1+\frac{H^{\mathcal{S}}_{n}}{H_{\text{ub}}}\right)\,.\end{split}$ (12)
-This reformulation expresses $w_{\text{EA},n}$ as a product of the $H^{\mathcal{T}}_{n}$ and a scaling factor that depends on the $H^{\mathcal{S}}_{n}/H_{\text{ub}}$.
-In this way, the teacher’s assessment of sample value is adaptively adjusted based on the student’s perspective.
+This reformulation expresses $w_{\text{EA},n}$ as the product of $H^{\mathcal{T}}_{n}$ and a scaling factor that depends on the $H^{\mathcal{S}}_{n}$.
+Thus, $w_{\text{EA},n}$ can be regarded as the teacher’s assessment of sample value, adaptively adjusted based on the student’s perspective.
 Depending on the entropy values, $w_{\text{EA}}$ behaves as follows:
@@
 $w_{\text{EA},n}=$ $\frac{H_{\text{ub}}}{2}$ if $H^{\mathcal{T}}_{n}\rightarrow H_{\text{ub}}\land H^{\mathcal{S}}_{n}\rightarrow 0$ (13b)
-$w_{\text{EA},n}=$ 00 0 if $H^{\mathcal{T}}_{n}\rightarrow 0\quad\forall H^{\mathcal{S}}_{n}$ (13c)
+$w_{\text{EA},n}=$ $0$ if $H^{\mathcal{T}}_{n}\rightarrow 0\quad\forall H^{\mathcal{S}}_{n}$ (13c)
 $w_{\text{EA},n}=$ $w_{\text{EA},n}$ otherwise.
@@
 Accuracy (%) of EA-methods vs.
-originals across teacher-student pairs, with relative improvements highlighted.
+baselines across teacher-student pairs, with relative improvements highlighted.
 Avg.
@@
 Best results are bolded, and second-best are underlined.
-Type Teacher ResNet32$\times$4 WRN-28-4 WRN-40-2 VGG13 VGG13 ResNet50 ResNet32$\times$4
+[…]
 79.42 78.60 75.61 74.64 74.64 79.34 79.42
-Student ResNet8$\times$4 WRN-16-2 WRN-40-1 VGG8 MN-V2 MN-V2 SN-V2
+[…]
 72.50 73.26 71.98 70.36 64.60 64.60 71.82 Avg. $\Delta$
@@
 MLD+LS [38] 78.28 77.20 75.56 75.22 70.94 71.19 78.76
-EA-MLD+LS 78.38 +0.09 77.60 +0.39 75.78 +0.22 75.38 +0.16 70.67 -0.27 71.36 +0.17 79.13 +0.37 +0.16+
+EA-MLD+LS 78.38 +0.10 77.60 +0.39 75.78 +0.22 75.38 +0.16 70.67 -0.27 71.36 +0.17 79.13 +0.37 +0.16+
 Feature ReviewKD [2] 75.63 76.39 74.45 74.45 70.37 69.89 77.78
@@
 For instance, when integrated with vanilla KD, the loss is defined as:
-$\begin{split}L_{\text{EA-KD}}&=\sum_{n=1}^{N}w_{\text{EA},n}\cdot L_{\text{KD},n}\\ &=\sum_{n=1}^{N}w_{\text{EA},n}\cdot\text{KL}(p^{\mathcal{T}}_{n}\|p^{\mathcal{S}}_{n})\,,\end{split}$ (14)
-where $L_{\text{KD},n}$ is the distillation loss for sample $x_{n}$, and $p^{\mathcal{T}}_{n}$ and $p^{\mathcal{S}}_{n}$ denote the teacher and student’s softened probability for $x_{n}$, respectively.
-Consequently, EA-KD introduces a more nuanced and dynamic knowledge transfer to standard KD frameworks, ensuring that valuable samples receive enhanced focus throughout training.
+$\begin{split}L_{\text{EA-KD}}&=\sum_{n=1}^{N}w_{\text{EA},n}\cdot L_{\text{KD},n},\end{split}$ (14)
+where $L_{\text{KD},n}$ is the distillation loss for sample $x_{n}$.
+As a result, EA-KD enhances standard KD methods by facilitating a more nuanced and adaptive transfer of knowledge, ensuring that informative samples receive increased focus throughout training.
 4 Experiments
@@
 Datasets.
-We used CIFAR-100 [12] (50k training, 10k validation images; 100 classes) and Tiny-ImageNet [14] (100k training, 5k validation images; 200 classes) to evaluate our method for image classification, and used MS-COCO [18] (118k training, 5k validation images; 80 classes) for object detection.
+We used CIFAR-100 [16] (50k training, 10k validation images; 100 classes), Tiny-ImageNet [18] (100k training, 5k validation images; 200 classes), and ImageNet [6] (1.28M training, 50k validation images; 1,000 classes) to evaluate our method for image classification.
+For object detection, we adopted MS-COCO [25] (118k training, 5k validation images; 80 classes).
 Furthermore, for LLM distillation, we employed five instruction-following datasets following [9]: Dolly [5], SInst [42], Vicuna [4], S-NI [43], and UnNI [12].
-Table 3: Results on Tiny-ImageNet.
-Top-1 (first row) and top-5 accuracy (second row) for both EA-methods and their original versions using ResNet32$\times$4 teacher and ResNet8$\times$4 student are reported.
-Teacher Student KD EA-KD MLD EA-MLD MLD+LS EA-MLD+LS FCFD EA-FCFD Avg. $\Delta$
-64.41 55.25 56.00 59.39 +3.39 61.91 62.65 +0.74 56.45 61.93 +5.48 60.12 60.51 +0.39 +2.50
-85.07 79.62 79.64 82.40 +2.76 83.77 84.47 +0.70 81.15 84.35 +3.20 82.90 83.22 +0.32 +1.74
-Table 4: Results on Tiny-ImageNet with Transformers. Accuracy (%) for EA-KD and KD using transformer-based teachers and a ResNet8$\times$4 student. Each result is averaged over three runs.
+[…]
+Teacher Student KD EA-KD MLD EA-MLD MLD+LS EA-MLD+LS FCFD EA-FCFD
+64.41 55.25 56.00 59.39 +3.39 61.91 62.65 +0.74 61.36 62.41 +1.05 60.12 60.51 +0.39
+Table 4: Results on ImageNet. Accuracy (%) of the ResNet34 teacher and ResNet18 student, averaged over three runs.
+Teacher Student KD [11] EA-KD KD+LS [38] DKD [47] EA-DKD DKD+LS [38] EA-DKD+LS PAD [46]
+73.31 69.75 71.03 71.79 +0.76 71.42 71.70 71.96 +0.26 71.88 71.99 +0.11 71.71
+[…]
 Teacher Student KD EA-KD
-ViT-B 71.12 ResNet8$\times$4 55.25 54.70 57.53 +2.83
+[…]
 DeiT-B 85.55 56.30 58.36 +2.06
 Swin-B 86.30 56.15 59.58 +3.43
-Table 5: Results on MS-COCO. Performance of EA-KD and KD using Faster-RCNN with FPN. Metrics include AP (overall), AP50, and AP75.
+Table 6: Results on MS-COCO. AP (overall), AP50, and AP75 are reported using Faster R-CNN with FPN.
 R-101 & R-18 R-50 & MV2
@@
 For fair comparisons, no hyperparameter was tuned for the EA-variants, except for the KD weight of MLD+LS [38] was reduced from 9.0 to 4.0 to avoid over-penalizing.
-Implementation Details.
-The EA-KD and other variants were evaluated on various CNNs (VGG [29], ResNet [8], WideResNet [36], MobileNet [27], ShuffleNet [23]), vision transformer teachers (ViT [6], DeiT [32], Swin [21]), and GPT2 [25] for LLM distillation.
-We used the training settings of [31] for vanilla KD and the respective settings of each KD framework for CIFAR-100 and Tiny-ImageNet, [1] for MS-COCO, and [37] for LLMs.
-We set $T^{\prime}=3$ based on our ablation study, except $T^{\prime}=2$ for transformer teachers.
-All results averaged over five runs, and analyses were conducted using a ResNet32$\times$4 teacher and a ResNet8$\times$4 student on CIFAR-100.
+[…]
 4.2 Results
@@
 These findings highlight EA-KD’s broad applicability and effectiveness in enhancing knowledge transfer across diverse types of KD frameworks.
-Tiny-ImageNet.
+Tiny-ImageNet and ImageNet.
 Tab.
-3 presents results on Tiny-ImageNet across SOTA frameworks enhanced with EA.
-For instance, EA-KD significantly improves top-1 accuracy by 3.39% and top-5 accuracy by 2.76% over standard KD, achieving performance close to FCFD with less than a 1% difference.
-These results highlight EA’s effectiveness on a larger and more class-diverse dataset.
-Further, Tab.
-4 highlights EA-KD’s ability to outperform KD when distilling from transformer-based teachers to a CNN student.
-While standard KD stopped improving the student when Swin-B was used as the teacher, EA-KD continued to enhance student performance, mitigating the capacity gap by allowing better teachers to produce better students.
+3 and Tab.
+4 show that EA-methods consistently outperform their baselines on both Tiny-Imagenet and ImageNet.
+On Tiny-ImageNet, EA-KD improves KD by 3.39% and achieves performance comparable to the novel FCFD [26].
+On ImageNet, EA-KD surpasses other adaptive KDs, including logit-level adaptive LS [38] and sample-level adaptive approach PAD [46].
+These results highlight EA-methods’ scalability on larger, more diverse datasets.
+Furthermore, Tab.
+5 underscores EA-KD’s advantage in distilling valuable knowledge from vision transformer-based teachers to CNN students.
+Unlike standard KD, which plateaued when using DeiT-B as the teacher, EA-KD consistently improved student performance, mitigating the capacity gap and utilizing the guidance of stronger teachers more effectively.
 MS-COCO.
@@
 This highlights EA-KD’s ability to effectively guide the student with valuable knowledge in a more complex visual task, thereby enhancing its performance.
-Table 6: LLM Distillation Results. Rouge-L [17] scores averaged over five seeds for KD, EA-KD, RKLD, and JSD on each dataset. SFT denotes a student fine-tuned on the dataset.
+Table 7: LLM Distillation Results. Rouge-L [24] scores averaged over five seeds on each dataset are reported. SFT denotes a student supervised fine-tuned on the dataset.
 Dataset
@@
 As shown in Tab.
-6, EA-KD consistently outperforms standard KD across datasets.
+7, EA-KD consistently outperforms standard KD across datasets.
 Moreover, EA-KD also surpasses two recently emerging methods in this field, Reverse KLD (RKLD) and Jensen-Shannon divergence (JSD), which aim to address KLD’s limitation of forcing the student to cover all modes of the complex LLM teacher distribution [9].
@@
 Finally, we highlight EA-methods computational efficiency across diverse KD frameworks.
-(a) KLD for Top and Bottom 10% Entropy Samples
-(b) Raw KLD and Training Accuracy (%)
-Figure 5: Comparison of KD and EA-KD During Training.
-EA-KD enhances focus on high entropy samples, leading to a more effective and efficient learning process, as reflected in the overall raw KLD and training accuracy.
 Ablation Study.
@@
 (i) Tab.
-7 shows that a $T^{\prime}$ of 3 consistently delivers optimal performance across model combinations, highlighting its robustness in reflecting sample value during distillation.
+8 shows that a $T^{\prime}$ of 3 consistently delivers optimal performance across model combinations, highlighting its robustness in reflecting sample value.
 (ii) As shown in Tab.
-8, the single-sided factor $w_{\text{base}}$ notably boosts baseline performance by emphasizing valuable samples, while $w_{\text{interact}}$ integrates teacher-student dynamics but may overly rely on the student responses, leading to limited gains.
-By combining these strengths, $w_{\text{EA}}$ achieves superior improvement by prioritizing valuable samples and adapting to the student’s learning dynamics, enabling a precise and tailored focus.
+9, $w_{\text{base}}$ enhances KD by emphasizing valuable samples under the teacher’s guidance.
+While $w_{\text{interact}}$ integrates student dynamics, it may introduce noise and dilute the teacher’s guidance, reducing effectiveness.
+However, when integrated into $w_{\text{EA}}$, $H^{\mathcal{S}}$ acts as a scaling factor for $H^{\mathcal{T}}$ (Eq.
+12).
+This ensures early training, where $H^{\mathcal{S}}$ is mostly high and the scaling near 1, is mainly guided by $H^{\mathcal{T}}$.
+As training progresses, $H^{\mathcal{S}}$ adaptively adjusts and tailors the weighting, leading to superior performance.
 Table 8: Impact of $T^{\prime}$. Setting $T^{\prime}$ to 3 yields the best results across architectures, demonstrating its robustness and consistency.
@@
 2 3 4
-ResNet32$\times$4 ResNet8$\times$4 73.90 75.46 75.22
+[…]
 WRN-28-4 WRN-16-2 74.89 75.79 75.62
-ResNet32$\times$4 SN-V2 75.51 75.91 75.72
+[…]
 Table 9: Impact of Reweighting Factors.
@@
 $w_{\text{base}}$ $w_{\text{interact}}$ $w_{\text{EA}}$
-KD [9] 73.33 $\uparrow$ 75.14 $\uparrow$ $\uparrow$ 74.76 $\uparrow$ $\uparrow$ 75.27 $\uparrow$
+KD [11] 73.33 $\uparrow$ 75.14 $\uparrow$ $\uparrow$ 74.76 $\uparrow$ $\uparrow$ 75.46 $\uparrow$
 MLD [14] 77.08 $\uparrow$ 77.47 $\uparrow$ $\uparrow$ 77.45 $\uparrow$ $\uparrow$ 77.65 $\uparrow$
 MLD+LS [38] 78.28 $\uparrow$ 78.30 $\uparrow$ $\uparrow$ 78.20 $\downarrow$ $\uparrow$ 78.38 $\uparrow$
-FCFD [19] 76.62 $\uparrow$ 77.37 $\uparrow$ $\uparrow$ 77.42 $\uparrow$ $\uparrow$ 77.44 $\uparrow$
-Refined Knowledge Focus and Alignment.
-Fig.
-5(a) illustrates the KLD for the top and bottom 10% entropy samples throughout training for both KD and EA-KD, with Fig.
-5(b) depicting the overall raw KLD and training accuracy.
-Notably, vanilla KD’s uniform distillation causes low-entropy samples to dominate the distillation loss, failing to transfer the more critical knowledge from higher-entropy samples, as discussed in Sec.
-3.
-Fig.
-2(a) further illustrates this, where KD shows increasingly larger accuracy gaps with the teacher on higher entropy samples.
-In contrast, EA-KD refined the student’s focus on high-entropy samples and efficiently reduced the raw KLD (Fig.
-5), achieving improved mimicry and generalization with more consistent alignment to the teacher across entropy segments (Fig.
+FCFD [26] 76.62 $\uparrow$ 77.50 $\uparrow$ $\uparrow$ 77.42 $\uparrow$ $\uparrow$ 77.44 $\uparrow$
+Distillation Loss and Sample Value.
+In Fig.
+5, we analyze the distribution of distillation loss across samples grouped by low- to high-entropy quartiles (Q1–Q4) for KD and EA-KD on CIFAR-100 and Tiny-ImageNet.
+Notably, KD loss is dominated by low-value samples throughout training on both datasets.
+The overshadowing effect in KD’s uniform distillation scheme hinders the transfer of critical knowledge from high-value samples, as discussed in Sec.
+3, ultimately leading to more significant accuracy gaps (Fig.
+2(a)).
+In contrast, EA-KD emphasizes the focus on valuable samples (Fig.
+5), bringing holistic performance improvements across entropy segments (Fig.
 2(a)).
@@
 Fig.
-6 compares the t-SNE visualizations [33] of KD and EA-KD with the high $H^{\mathcal{S}}$ and $w_{\text{EA}}$ samples highlighted.
+6 compares the t-SNE visualizations [41] of KD and EA-KD with the high $H^{\mathcal{S}}$ and $w_{\text{EA}}$ samples highlighted [^4] See Sec.
+B.1 for further t-SNE visualizations showing the student’s evolving sample focus over training epochs..
 Similar to Fig.
@@
 This suggests that, despite having a different view from the teacher (Fig.
-3), the student’s evolving capacity enables $H^{\mathcal{S}}$ to capture the samples crucial for its own learning progress [^3] See Sec.
-A.1 for further t-SNE visualizations showing the student’s evolving sample focus over training epochs..
-In EA-KD, integrating both $H^{\mathcal{T}}$ and $H^{\mathcal{S}}$ in $w_{\text{EA}}$ allows us to harness the complementary strengths of the teacher’s informed assessment and the student’s dynamic learning needs.
-Consequently, EA-KD achieves superior class separability as indicated by a higher Calinski-Harabasz (CH) index.
+3), the student’s evolving capacity enables $H^{\mathcal{S}}$ to capture the samples crucial for its own learning progress.
+In EA-KD, integrating both $H^{\mathcal{T}}$ and $H^{\mathcal{S}}$ in $w_{\text{EA}}$ captures both the teacher’s informed assessment and the student’s dynamic learning needs.
+As a result, EA-KD achieves superior class separability as reflected by a higher Calinski-Harabasz (CH) index.
+Figure 5: Loss Distribution vs. Entropy Quartiles over Epochs. Low-value samples (Q1, Q2; blue) dominate the KD loss, whereas EA-KD places more focus on high-value ones (Q3, Q4; red).
+Figure 6: t-SNE of Students.
+Top 10% high $H^{\mathcal{S}}$ samples for KD (left) and $w_{\text{EA}}$ samples for EA-KD (right) are highlighted with triangles.
+High $H^{\mathcal{S}}$ samples cluster near decision boundaries, underscoring their learning value for the student.
 Consistency in Entropy Levels.
@@
 As a result, the student more effectively mimics the teacher’s response and achieves improved performance.
-Figure 6: t-SNE of Students.
-Top 10% high $H^{\mathcal{S}}$ samples for KD (left) and $w_{\text{EA}}$ samples for EA-KD (right) are highlighted with triangles.
-High $H^{\mathcal{S}}$ samples cluster near decision boundaries, underscoring their learning value for the student.
-EA-KD achieves better class separability as indicated by a higher CH index.
 Figure 7: Loss Surface and Differences in Area@1.6 for DKD and EA-DKD Students Across Epochs.
@@
 7, lower right) underscores EA-DKD’s closer alignment with the teacher’s surface throughout training.
-The enhanced generalization in EA-DKD also extends to the robustness of its hyperparameter.
-DKD was introduced to decouple target class KD (TCKD) and non-target class KD (NCKD) in vanilla KD, balancing them with hyperparameters $\alpha$ and $\beta$ [38].
+The enhanced generalization in EA-DKD also extends to hyperparameter robustness.
+DKD was introduced to decouple target class KD (TCKD) and non-target class KD (NCKD) in vanilla KD, balancing them with hyperparameters $\alpha$ and $\beta$.
 While $\alpha$ remains stable around 1.0, adjusting $\beta$ from 1.0 to 10.0 leads to fluctuating performance [47], with a variance of 0.31 (see Fig.
@@
 Notably, EA-DKD effectively reduces this fluctuation to a variance of 0.10 and consistently improves the performance.
-Overall, the enhanced generalizability and robustness may be attributed to the synergy between DKD and EA in handling different aspects of knowledge transfer: DKD prevents NCKD from being overshadowed by TCKD at the class level, while EA ensures valuable samples—often rich in NCKD—are not dominated by simpler ones at the sample level.
+This enhanced robustness may be attributed to the synergy between DKD and EA in handling different aspects of knowledge transfer: DKD prevents NCKD from being overshadowed by TCKD at the class level, while EA ensures valuable samples—often rich in NCKD—are not dominated by simpler ones at the sample level.
 Additionally, the dynamic focus in EA compensates for the static nature of $\beta$.
-Together, EA-DKD facilitates a more balanced and efficient knowledge transfer process, both class-wise and sample-wise.
+Together, EA-DKD facilitates a nuanced and balanced knowledge transfer process, both class-wise and sample-wise.
 Computational Efficiency.
 Fig.
-1 illustrates the efficiency of EA-methods, showing significant performance improvements on various SOTA KD frameworks with negligible cost.
-This remarkable computational efficiency, combined with its streamlined integration into both logit- and feature-based KD frameworks, suggests the potential of our method as a versatile and accessible enhancement for KD.
+1 illustrates the efficiency of EA-methods, showing significant performance improvements across various SOTA KD frameworks with negligible cost.
+This remarkable efficiency, combined with its streamlined integration into both logit- and feature-based KD methods, underscores the potential of our method as a versatile and practical enhancement for KD.
 5 Conclusion
-In this paper, we examined existing KD methods from a novel perspective, revealing the limitation of their inherent uniform distillation scheme, which often hinders the transfer of valuable, high-entropy samples that contain critical knowledge.
-To overcome this, we introduced EA-KD, a plug-and-play KD approach that dynamically reweights the distillation loss, directing the learning process toward valuable samples.
-EA-KD consistently improves SOTA KD methods across image classification, object detection, and LLM distillation, all with negligible cost.
-We believe EA-KD showcases a great paradigm of the meticulous handling of knowledge transfer, offering sample-wise adaptability while integrating the student’s evolving learning dynamics into the distillation process.
+In this paper, we revisited existing KD methods from a novel perspective and revealed a key limitation in their inherent uniform distillation strategy, which often hinders the transfer of high-entropy samples that carry critical knowledge.
+To address this, we proposed EA-KD, a plug-and-play KD approach that dynamically reweights the distillation loss, directing the learning focus toward valuable samples.
+EA-KD consistently enhances representative KD baselines across image classification, object detection, and LLM distillation, all with negligible cost.
+We believe EA-KD showcases a great paradigm of the meticulous handling of knowledge transfer, adapting KD to the varying learning value of samples while accounting for the student’s evolving learning dynamics throughout the distillation process.
 Supplementary Material
-Appendix A Additional Analysis and Visualizations
-This document provides additional analysis to complement the main paper.
+Appendix A Additional Results and Comparison
+A.1 Comparison with TTM and WTTM
+TTM [48] removes student temperature in KD, revealing an inherent Rényi entropy regularizer that could improve generalization.
+WTTM further up-weights uncertain samples using the power sum of the teacher’s output, akin to our $w_{\text{base}}$ [^5] In fact, they also noted teacher entropy as a potential weighting factor, but left its systematic exploration for future work..
+Notably, TTM’s Rényi entropy is an inherent effect of its structural modification, while EA-KD actively uses Shannon entropy to prioritize valuable samples.
+Moreover, unlike WTTM’s static teacher-based weighting, EA-KD dynamically adapts to the student’s learning (via $H^{\mathcal{S}}$), yielding stronger enhancement to TTM (Tab.
+A1).
+Finally, while their methods combine with feature-based KDs by adding and balancing their losses, EA-KD offers direct integration to both logit- and feature-based KDs through reweighting.
+Table A1: TTM, WTTM, and EA-TTM on CIFAR-100 over 5 runs. Unlike WTTM, EA-TTM consistently improves TTM, highlighting the superiority of dynamic over static weighting.
+Teacher Student TTM WTTM EA-TTM (Ours)
+[…]
+[…]
+[…]
+A.2 MS-COCO
+Tab.
+A2 reports additional results for MS-COCO, where EA-DKD is compared against DKD [47] using a ResNet-101 to ResNet-18 distillation setting.
+The results show that EA-DKD consistently improves upon DKD across all evaluation metrics.
+Table A2: More Results on MS-COCO.
+[…]
+DKD [39] 35.05 56.60 37.54
+EA-DKD 35.16+0.11 56.75 +0.15 37.82 +0.28
+A.3 LLM Distillation
+Tab.
+A3 extends the comparison of EA-KD with Skewed KLD (SKLD) and SRKLD from the recent DistiLLM [15] method, under the same off-policy setting without pre-training corpus.
+EA-KD outperforms these stronger methods on most datasets, highlighting its overall effectiveness.
+Table A3: More LLM Distillation Results.
+Method Dolly SInst Vicuna S-NI UnNI Avg.
+SKLD (DistiLLM) 24.03 10.66 14.70 17.99 21.18 17.71
+SRKLD (DistiLLM) 24.48 10.35 14.88 16.53 19.68 17.19
+EA-KD 24.95 10.59 16.41 18.27 21.46 18.34
+Appendix B Additional Analysis and Visualizations
+This section provides additional analysis to complement the main paper.
 It includes a detailed examination of high student entropy samples, loss landscape comparisons for KD and EA-KD, and t-SNE visualizations for various KD frameworks and their EA-enhanced variants.
-A.1 Analysis of High Student Entropy Samples
+B.1 Analysis of High Student Entropy Samples
 Complementary to the teacher entropy analysis in Sec.
@@
 As shown in Fig.
-1(a), high student entropy samples correlate with larger teacher-student accuracy gaps similar to the teacher entropy segments (Fig.
+1(a), these samples correlate with larger teacher-student accuracy gaps similar to the teacher entropy segments (Fig.
 2(a)).
 Additionally, Fig.
-1(b) presents the students’ t-SNE visualizations across training epochs for KD and EA-KD, with high-entropy samples ($H^{\mathcal{S}}$) and reweighted samples ($w_{\text{EA}}$) highlighted, respectively.
-Notably, these high-entropy samples also cluster near decision boundaries similar to the teacher’s t-SNE (Fig.
+1(b) presents the students’ t-SNE visualizations across training epochs for KD and EA-KD, with high $H^{\mathcal{S}}$ and high $w_{\text{EA}}$ samples highlighted, respectively.
+Notably, these samples also cluster near decision boundaries similar to the teacher’s t-SNE (Fig.
 2(b)).
@@
 (b) t-SNE of Students Across Training Epochs.
-Figure A1: High Student Entropy Analysis in KD and EA-KD.
+Figure B1: High Student Entropy Analysis in KD and EA-KD.
 (a) Higher student entropy samples correlate with larger accuracy gaps in KD.
@@
 EA-KD adapts to this dynamic, achieving enhanced class separability over epochs.
-A.2 Loss Landscape Analysis for KD and EA-KD
+Figure B2: Loss Surface and Differences in Area@1.6 for KD and EA-KD Students Across Epochs.
+Contour plots illustrate the loss landscapes of KD (first row) and EA-KD (second row) across training epochs.
+The line plot (lower right) tracks the differences in Area@1.6 between the teacher and students over epochs.
+EA-KD exhibits a more stable and robust loss surface, with greater Area@1.6 earlier in training, signifying a more efficient learning process.
+B.2 Loss Landscape Analysis for KD and EA-KD
 We further analyze the robustness of KD and EA-KD by examining their loss landscapes across training epochs.
 As shown in Fig.
-A2, EA-KD achieves a consistently smoother loss surface compared to KD (e.g.
+B2, EA-KD achieves a consistently smoother loss surface compared to KD (e.g.
 $\sigma^{2}$ of 217.68 vs.
@@
 In contrast, EA-KD achieves values of 0.01 and 0.05, respectively, underscoring its improved learning efficiency by enabling the student to assimilate the key knowledge.
-These results highlight EA-KD’s ability to ensure a more efficient and effective optimization, facilitating enhanced generalizability compared to standard KD.
-A.3 t-SNE of KD Frameworks and EA-methods
+These results highlight EA-KD’s ability to ensure a more efficient optimization, facilitating enhanced generalizability compared to standard KD.
+B.3 t-SNE of KD Frameworks and EA-methods
 The t-SNE visualizations of students from various KD frameworks and their EA-method variants, alongside the teacher’s representation, are shown in Fig.
-A3.
+B3.
 EA-methods consistently exhibit more distinct and well-defined class clusters compared to their respective baselines, as evidenced by higher CH indices.
@@
 These findings highlight EA-KD’s versatility in enhancing class separability and improving knowledge transfer across diverse KD frameworks.
-Figure A2: Loss Surface and Differences in Area@1.6 for KD and EA-KD Students Across Epochs.
-Contour plots illustrate the loss landscapes of KD (first row) and EA-KD (second row) across training epochs.
-The line plot (lower right) tracks the differences in Area@1.6 between the teacher and students over epochs.
-EA-KD exhibits a more stable and robust loss surface, with greater Area@1.6 earlier in training, signifying a more efficient learning process.
-Figure A3: t-SNE Visualizations of EA-methods vs.
+Figure B3: t-SNE Visualizations of EA-methods vs.
 Baselines.
@@
 EA-methods consistently achieve higher CH indices, indicating better class separability.
+Appendix C Statistics for Table 2 and 4
+Tab.
+C4 and Tab.
+C5 present the mean and standard deviation of EA-methods on CIFAR-100 and ImageNet across multiple runs, respectively.
+Since baseline papers typically report only mean values, formal statistical tests such as p-value computation are not applicable.
+Nonetheless, EA-methods consistently demonstrate improved performance with low variance, highlighting the stable and reliable gains.
+Table C4: Results on CIFAR-100. Mean accuracy (%) and standard deviation across five runs are reported.
+[…]
+79.42 78.60 75.61 74.64 74.64 79.34 79.42
+[…]
+72.50 73.26 71.98 70.36 64.60 64.60 71.82
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table C5: Results on ImageNet. Mean accuracy (%) and standard deviation across three runs are reported for EA-methods.
+Teacher Student KD [10] EA-KD KD+LS [30] DKD [39] EA-DKD DKD+LS [30] EA-DKD+LS PAD [38]
+[…]